will you answer the following question |x+2|-|x-1|

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
In this type of question break these at point of discontinuity
i.e. at 1 & -2
 
When x
 
-x -2 + x -1 = 0
-3 = 0 it is unacceptable hence x can't be less than -2
 
Now if -2
x+2 + x -1 = 0
2x = -1
x = -1/2
 
If x>1
 
x +2 -x +1 = 0
3 = 0 hence x can't be greater than 1
 
Bir
5 years ago
|x|=x:x>0......|x|=-x:x